OpenStudy (anonymous):

a rectangular solid whose edges are 4 cm, 5 cm, and 7cm is inscribed in a sphere., Find the volume and the surface area of the sphere.

OpenStudy (anonymous):

Assuming length of solid=7 cm Width=5 cm Diagonal of base=sqrt(7 sqre+5sqre)=sqrt(49+25)=sqrt74 Height=4 cm Diagonal of solid=sqrt(sqre(sqrt74)+4 sqre)=sqrt(74+16)=sqrt90 Diagonal of sphere d=sqrt90=3 sqrt 10 Radius of sphere r=(3 sqrt 10 )/2 Volume of sphere V =4/3( pi r cube )=? Surface area of sphere A=4 pi r sqre )=?

OpenStudy (anonymous):

i think u are missing one thing